# Tight concentration inequalities for quantum adversarial setups exploiting permutation symmetry

### v2

@@ II.2 Generalization to the case with symmetry restrictions
where $Q_{\bm{j}}Q_{\bm{j}^{\prime}}=\delta_{\bm{j}\bm{j}^{\prime}}$. Then, the $\pi^{\otimes n}(G^{\times n})$-invariant state ${\rho}_{A^{n}E}^{G}$ is decomposed as
@@
$s[\bm{m}](\rho)$ $\coloneqq\!\sum_{\bm{j}\in{\cal T}(\bm{m})}\!(V_{\sigma_{\bm{j}}}^{D})^{\dagger}\mathrm{Tr}_{{\cal H}_{\bm{j}}^{R}}\otimes\mathrm{Id}_{{\cal H}_{\bm{j}}^{D}}\!\bigl{(}Q_{\bm{j}}\rho Q_{\bm{j}}|_{{\cal H}_{\bm{j}}^{R}\otimes{\cal H}_{\bm{j}}^{D}}\bigr{)}V_{\sigma_{\bm{j}}}^{D}.$ (100)
@@
$\rho_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}E}=\!\!\bigoplus_{\begin{subarray}{c}(\bm{n}_{1},\ldots,\bm{n}_{k})\end{subarray}}\!\!P^{D}_{\bm{j}_{0}(\bm{m});\bm{n}_{1},\ldots,\bm{n}_{k}}\rho_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}E}P^{D}_{\bm{j}_{0}(\bm{m});\bm{n}_{1},\ldots,\bm{n}_{k}}.$ (106)
@@
(ii) For any $\rho\in{\cal D}({\cal H}_{\bm{j}_{0}(\bm{m})}^{D}\otimes{\cal H}_{E})$ that is $S_{n_{1}}\times\cdots\times S_{n_{k}}$-invariant on ${\cal H}_{\bm{j}_{0}(\bm{m})}^{D}$, we have
@@
(iii) For any $\rho\in{\cal D}({\cal H}_{\bm{j}_{0}(\bm{m})}^{D})$ that is $S_{n_{1}}\times\cdots\times S_{n_{k}}$-invariant and for any $\bm{n}_{j}\in Y_{n_{j}}^{d_{j}}$ for $j=1,\ldots,k$, we have
$\begin{split}&\mathrm{supp}\bigl{(}\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}(P^{D}_{\bm{j}_{0}(\bm{m});\bm{n}_{1},\ldots,\bm{n}_{k}}\rho P^{D}_{\bm{j}_{0}(\bm{m});\bm{n}_{1},\ldots,\bm{n}_{k}})\bigr{)}\\ &\leq\prod_{j=1}^{k}f_{q}(n_{j},d_{j};\bm{n}_{j})\\ &\quad\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\!\biggl{(}\bigotimes_{j=1}^{k}\mathbb{E}_{U_{j}^{D}\sim\mathrm{Haar}(d_{j})}\Bigl{[}\Bigl{(}U_{j}^{D}\mathrm{diag}\Bigl{(}\frac{\bm{n}_{j}}{n_{j}}\Bigr{)}(U_{j}^{D})^{\dagger}\Bigr{)}_{{\cal H}_{j}^{D}}^{\otimes n_{j}}\Bigr{]}\biggr{)}.\end{split}$ (109)
@@
$\begin{split}{\cal R}_{A^{n}}^{G}(\rho)&\coloneqq\bigoplus_{\bm{m}\in Z_{n}^{k}}r[\bm{m}]\circ\Bigl{(}\mathrm{Tr}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{R}}\otimes\tilde{\cal R}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\Bigr{)}\\ &\qquad\quad\biggl{(}\sum_{\bm{j}\in{\cal T}(\bm{m})}V_{\sigma_{\bm{j}}}^{\dagger}Q_{\bm{j}}\rho Q_{\bm{j}}V_{\sigma_{\bm{j}}}\biggl{|}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{R}\otimes{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\biggr{)},\end{split}$ (111)
@@
$\rho_{A^{n}E}^{G}=\bigoplus_{\bm{m}\in Z_{n}^{k}}\bigoplus_{\begin{subarray}{c}(\bm{n}_{1},\ldots,\bm{n}_{k})\\ \in Y_{n_{1}}^{d_{1}}\times\cdots\times Y_{n_{k}}^{d_{k}}\end{subarray}}P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}}\rho_{A^{n}E}^{G}P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}},$ (116)
@@
$\begin{split}&\mathrm{supp}\Bigl{(}\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\bigl{(}s[\bm{m}](P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}}\rho_{A^{n}E}^{G}P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}})\bigr{)}\Bigr{)}\\ &\leq\prod_{j=1}^{k}f_{q}(n_{j},d_{j};\bm{n}_{j})\\ &\quad\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\!\biggl{(}\bigotimes_{j=1}^{k}\mathbb{E}_{U_{j}^{D}\sim\mathrm{Haar}(d_{j})}\Bigl{[}\Bigl{(}U_{j}^{D}\mathrm{diag}\Bigl{(}\frac{\bm{n}_{j}}{n_{j}}\Bigr{)}(U_{j}^{D})^{\dagger}\Bigr{)}_{{\cal H}_{j}^{D}}^{\otimes n_{j}}\Bigr{]}\biggr{)}.\end{split}$ (118)
@@
$\begin{split}&s[\bm{m}]\bigl{(}\tau_{\bm{U}}(\bm{n}_{1},\ldots,\bm{n}_{k})\bigr{)}\\ &=\mathrm{Mult}_{\bm{m}/n}(\bm{m})\bigotimes_{j=1}^{k}\Bigl{(}U_{j}^{D}\mathrm{diag}\Bigl{(}\frac{\bm{n}_{j}}{n_{j}}\Bigr{)}(U_{j}^{D})^{\dagger}\Bigr{)}_{{\cal H}_{j}^{D}}^{\otimes n_{j}}.\end{split}$ (119)
@@
$\begin{split}&\mathrm{supp}\bigl{(}S_{A^{n}}^{G}(P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}}\rho_{A^{n}E}^{G}P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}})\bigr{)}\\ &\leq f_{c}(n,k;\bm{m})\prod_{j=1}^{k}f_{q}(n_{j},d_{j};\bm{n}_{j})\ket{\varnothing}\!\bra{\varnothing}_{{\cal H}_{\bm{j}_{0}}^{R}}\\ &\quad\otimes\mathbb{E}_{\bm{U}\sim\times_{j=1}^{k}\mathrm{Haar}(d_{j})}\Bigl{[}\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\circ s[\bm{m}]\Bigl{(}\tau_{\bm{U}}(\bm{n}_{1},\ldots,\bm{n}_{k})^{\otimes n}\Bigr{)}\Bigr{]}\end{split}$ (120)
@@
II.3 Concentration inequality for an independent and identical measurement on an adversarial state
@@
for any $\bm{i}\in{\cal X}^{\times n}$. What we want to know is a restriction imposed on the probability $\mathrm{P}_{\tilde{\rho}}(\bm{m})$ from the fact that the independent and identical measurements are performed on the permutation-invariant state. To know this, let $\mathcal{P}$ be the probability simplex, i.e., the set of probability vectors $\bm{p}=(p_{1},\ldots,p_{k})$, and $\mathcal{Q}\subseteq\mathcal{P}$ be its subset defined as
@@
Then, for a particular sequence $\bm{i}=(i_{1},\ldots,i_{n})$ of the measurement outcomes, a conditional probability of the $n$-th outcome $i_{n}$ conditioned on any sequence $\bm{i}^{\prime}=(i_{1},\ldots,i_{n-1})$ of the other $n-1$ outcomes should be in $\mathcal{Q}$ due to the definition of $\mathcal{Q}$. Let $\bm{m}^{\prime}=(n^{\prime}_{1},\ldots,n^{\prime}_{k})$ with $n^{\prime}_{j}=\chi^{\prime}_{j}(\bm{i}^{\prime})$, where $\bm{\chi}^{\prime}:{\cal X}^{\times(n-1)}\to Z_{n-1}^{k}$ is defined similarly to Eq. (33) for $(n-1)$ tuple, and let $\bm{m}^{\prime}[j]$ be defined as a $k$-tuple that adds one to $j$-th element of $\bm{m}^{\prime}$, i.e., $\bm{m}^{\prime}[j]=(n^{\prime}_{1},\ldots,n^{\prime}_{j-1},n^{\prime}_{j}+1,n^{\prime}_{j+1},\ldots,n^{\prime}_{k})$. By applying the above restriction on the conditional probability to Eq. (125), we have
$\begin{split}\left(\sum_{j=1}^{k}\frac{\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime}[j])}{|{\cal T}(\bm{m}^{\prime}[j])|}\right)^{-1}\!\!\left(\frac{\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime}[1])}{|{\cal T}(\bm{m}^{\prime}[1])|},\ldots,\frac{\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime}[k])}{|{\cal T}(\bm{m}^{\prime}[k])|}\right)\in\mathcal{Q}.\end{split}$ (127)
(Note, if only this condition is satisfied, the following argument holds for any classical probability mass function.) This restriction on $\mathrm{P}_{\tilde{\rho}}(\bm{m})$ for each $\bm{m}$ is strict, but is not very convenient for the later analysis. So, we consider a looser version for this that is easier to analyze. Let ${\cal R}\coloneqq\{(p_{1},\ldots,p_{k})\in{\cal P}:\sum_{j=1}^{k}a_{j}p_{j}\leq 0\}$ be a convex polytope with a nonempty interior such that ${\cal Q}\subseteq{\cal R}$ (see Fig. 1). Let $R$ be a boundary hyperplane $R\coloneqq\{(p_{1},\ldots,p_{k})\in{\cal P}:\sum_{j=1}^{k}a_{j}p_{j}=0\}$, which separates ${\cal P}$ by ${\cal R}$ and ${\cal P}\setminus{\cal R}$. This is a general form of a convex polytope and its boundary hyperplane that separates ${\cal P}$ into two since a hyperplane has the form $\sum_{i}a_{i}p_{i}=b$ in general and one can change $a_{i}\to a_{i}-b$ for all $i$ and $b\to 0$ exploiting the fact $\sum_{i}p_{i}=1$. In the following, we restrict our attention to the case where the complement ${\cal P}\setminus{\cal R}$ of ${\cal R}$ contains only one probability vector that corresponds to having a deterministic outcome, say $\bm{p}_{1}\coloneqq(1,0,\ldots,0)$. This may lead to a looser bound when applied to a specific problem or limit the applicability of the obtained bound, but as can be shown later, it still has a wide range of applications. The coefficients $\{a_{j}\}_{j=1}^{k}$ that defines the hyperplane $R$ in this case should satisfy $a_{1}>0$, $a_{j}\leq 0$ for $j=2,\ldots,k$. By redefining $b_{j}=-a_{j}/a_{1}(\geq 0)$ for $j=2,\ldots,k$, we have
${\cal R}$$\coloneqq\left\{(p_{1},\ldots,p_{k})\in{\cal P}:p_{1}-\biggl{(}\sum_{j=2}^{k}b_{j}p_{j}\biggr{)}\leq 0\right\},$ (128)
@@
where $\bm{u}\coloneqq(1,-b_{2},\ldots,-b_{k})$. Since ${\cal Q}\subseteq{\cal R}$ by assumption, we have
$\begin{split}\left(\sum_{j=1}^{k}\frac{\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime}[j])}{|{\cal T}(\bm{m}^{\prime}[j])|}\right)^{-1}\!\!\left(\frac{\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime}[1])}{|{\cal T}(\bm{m}^{\prime}[1])|},\ldots,\frac{\mathrm{P}_{\tilde{\rho}}(\bm{m}^{\prime}[k])}{|{\cal T}(\bm{m}^{\prime}[k])|}\right)\in\mathcal{R},\end{split}$ (130)
Then, with this constraint, we will find an upper bound on $\mathrm{P}_{\tilde{\rho}}(\bm{m})$ for an arbitrary frequency distribution $\bm{m}=(n_{1},\ldots,n_{k})$ of the outcomes. Let $\bm{l}=(l_{1},\ldots,l_{k})$ be a frequency distribution with $\sum_{j=1}^{k}l_{j}=n-1$. Let us further define that $\bm{l}[j]=(l_{1},\ldots,l_{j}+1,\ldots,l_{k})$ for $j=1,\ldots,k$. Now, for later use, we consider the following. Let $\tilde{\cal P}$ be the set of $k$-dimensional lattice points defined as
$\tilde{\cal P}\coloneqq\{(n_{1},\ldots,n_{k})/n\in{\cal P}:\forall j\in\{1,\ldots,k\},n_{j}\in\mathbb{Z}\}.$ (131)
Let $R^{\prime}$ be the hyperplane that translates $R$ by $\sqrt{2}/n$ towards $\bm{p}_{1}$. Then, define a sublattice $\tilde{\cal R}$ as the intersection of $\tilde{\cal P}$ and the region sandwiched by two hyperplanes $R$ and $R^{\prime}$ including $R$ and $R^{\prime}$ themselves. More explicitly, notice that $\bm{u}\in\mathbb{R}^{k}$ is the vector perpendicular to the hyperplane $R_{0}\coloneqq\{\bm{x}\in\mathbb{R}^{k}:\bm{u}\cdot\bm{x}=0\}$ towards $\bm{p}_{1}$, and that $\bm{v}\coloneqq(1,\ldots,1)/\sqrt{k}$ is a unit vector perpendicular to the hyperplane ${\cal P}$. Then, by defining
$\bm{w}\coloneqq\bm{u}-(\bm{v}\cdot\bm{u})\bm{v},$ (132)

### v3

@@ 2.2 Generalization to the case with symmetry restrictions
where $Q_{\bm{j}}Q_{\bm{j}^{\prime}}=\delta_{\bm{j}\bm{j}^{\prime}}Q_{\bm{j}}$. Then, the $\pi^{\otimes n}(G^{\times n})$-invariant state ${\rho}_{A^{n}E}^{G}$ is decomposed as
@@
$s[\bm{m}](\rho)$ $\coloneqq\!\sum_{\bm{j}\in{\cal T}(\bm{m})}\!(V_{\sigma_{\bm{j}}}^{D})^{\dagger}\mathrm{Tr}^{(\bm{j})}_{{\cal H}_{\bm{j}}^{R}}\bigl{(}Q_{\bm{j}}\rho Q_{\bm{j}}|_{{\cal H}_{\bm{j}}^{R}\otimes{\cal H}_{\bm{j}}^{D}}\bigr{)}V_{\sigma_{\bm{j}}}^{D}.$ (102)
@@
$\begin{split}&\rho_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}E}\\ &=\!\!\bigoplus_{\begin{subarray}{c}(\bm{n}_{1},\ldots,\bm{n}_{k})\end{subarray}}\!\!P^{D}_{\bm{j}_{0}(\bm{m});\bm{n}_{1},\ldots,\bm{n}_{k}}\rho_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}E}P^{D}_{\bm{j}_{0}(\bm{m});\bm{n}_{1},\ldots,\bm{n}_{k}}.\end{split}$ (108)
@@
(ii) For any $\rho\in{\cal D}({\cal H}_{\bm{j}_{0}(\bm{m})}^{D}\otimes{\cal H}_{E})$ that is $(S_{n_{1}}\times\cdots\times S_{n_{k}})$-invariant on ${\cal H}_{\bm{j}_{0}(\bm{m})}^{D}$, we have
@@
(iii) For any $\rho\in{\cal D}({\cal H}_{\bm{j}_{0}(\bm{m})}^{D})$ that is $(S_{n_{1}}\times\cdots\times S_{n_{k}})$-invariant and for any $\bm{n}_{j}\in Y_{n_{j}}^{d_{j}}$ for $j=1,\ldots,k$, we have
$\begin{split}&\mathrm{supp}\bigl{(}\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}(P^{D}_{\bm{j}_{0}(\bm{m});\bm{n}_{1},\ldots,\bm{n}_{k}}\rho P^{D}_{\bm{j}_{0}(\bm{m});\bm{n}_{1},\ldots,\bm{n}_{k}})\bigr{)}\\ &\leq\prod_{j=1}^{k}f_{q}(n_{j},d_{j};\bm{n}_{j})\,\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\!\biggl{(}\bigotimes_{j=1}^{k}\\ &\qquad\mathbb{E}_{U_{j}^{D}\sim\mathrm{Haar}(d_{j})}\Bigl{[}\Bigl{(}U_{j}^{D}\mathrm{diag}\Bigl{(}\frac{\bm{n}_{j}}{n_{j}}\Bigr{)}(U_{j}^{D})^{\dagger}\Bigr{)}_{{\cal H}_{j}^{D}}^{\otimes n_{j}}\Bigr{]}\biggr{)}.\end{split}$ (111)
@@
$\begin{split}{\cal R}_{A^{n}}^{G}(\rho)&\coloneqq\bigoplus_{\bm{m}\in Z_{n}^{k}}r[\bm{m}]\circ\tilde{\cal R}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\circ\mathrm{Tr}^{(\bm{j}_{0}(\bm{m}))}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{R}}\\ &\qquad\quad\biggl{(}\sum_{\bm{j}\in{\cal T}(\bm{m})}V_{\sigma_{\bm{j}}}^{\dagger}Q_{\bm{j}}\rho Q_{\bm{j}}V_{\sigma_{\bm{j}}}\biggl{|}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{R}\otimes{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\biggr{)},\end{split}$ (113)
@@
$\begin{split}&\rho_{A^{n}E}^{G}\\ &=\bigoplus_{\bm{m}\in Z_{n}^{k}}\bigoplus_{\begin{subarray}{c}(\bm{n}_{1},\ldots,\bm{n}_{k})\\ \in Y_{n_{1}}^{d_{1}}\times\cdots\times Y_{n_{k}}^{d_{k}}\end{subarray}}P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}}\rho_{A^{n}E}^{G}P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}},\end{split}$ (118)
@@
$\begin{split}&\mathrm{supp}\Bigl{(}\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\bigl{(}s[\bm{m}](P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}}\rho_{A^{n}E}^{G}P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}})\bigr{)}\Bigr{)}\\ &\leq\prod_{j=1}^{k}f_{q}(n_{j},d_{j};\bm{n}_{j})\,\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\biggl{(}\bigotimes_{j=1}^{k}\\ &\quad\mathbb{E}_{U_{j}^{D}\sim\mathrm{Haar}(d_{j})}\Bigl{[}\Bigl{(}U_{j}^{D}\mathrm{diag}\Bigl{(}\frac{\bm{n}_{j}}{n_{j}}\Bigr{)}(U_{j}^{D})^{\dagger}\Bigr{)}_{{\cal H}_{j}^{D}}^{\otimes n_{j}}\Bigr{]}\biggr{)}.\end{split}$ (120)
@@
$\begin{split}&s[\bm{m}]\bigl{(}\bigl{[}\tau_{\bm{U}}(\bm{n}_{1},\ldots,\bm{n}_{k})\bigr{]}^{\otimes n}\bigr{)}\\ &=\mathrm{Mult}_{\bm{m}/n}(\bm{m})\bigotimes_{j=1}^{k}\Bigl{(}U_{j}^{D}\mathrm{diag}\Bigl{(}\frac{\bm{n}_{j}}{n_{j}}\Bigr{)}(U_{j}^{D})^{\dagger}\Bigr{)}_{{\cal H}_{j}^{D}}^{\otimes n_{j}}.\end{split}$ (121)
@@
$\begin{split}&\mathrm{supp}\bigl{(}S_{A^{n}}^{G}(P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}}\rho_{A^{n}E}^{G}P_{\bm{m};\bm{n}_{1},\ldots,\bm{n}_{k}})\bigr{)}\\ &\leq f_{c}(n,k;\bm{m})\prod_{j=1}^{k}f_{q}(n_{j},d_{j};\bm{n}_{j})\ket{\varnothing}\!\bra{\varnothing}_{{\cal H}_{\bm{j}_{0}}^{R}}\\ &\quad\otimes\mathbb{E}_{\bm{U}\sim\times_{j=1}^{k}\mathrm{Haar}(d_{j})}\Bigl{[}\tilde{\cal S}_{{\cal H}_{\bm{j}_{0}(\bm{m})}^{D}}\circ s[\bm{m}]\Bigl{(}\\ &\hskip 85.35826pt\tau_{\bm{U}}(\bm{n}_{1},\ldots,\bm{n}_{k})^{\otimes n}\Bigr{)}\Bigr{]}\end{split}$ (122)
@@
2.3 Concentration inequality for an independent and identical measurement on an adversarial state
@@
for any $\bm{i}\in{\cal X}^{\times n}$. What we want to know is a restriction imposed on the probability $\mathrm{P}_{\tilde{\rho}}(\bm{m})$ from the fact that the independent and identical measurements are performed on the permutation-invariant state. To know this, let $\mathcal{P}$ be the probability simplex, i.e., the set of probability vectors $\bm{p}=(p_{1},\ldots,p_{k})$ with $p_{i}\geq 0$ and $\sum_{i=1}^{k}p_{i}=1$, and $\mathcal{Q}\subseteq\mathcal{P}$ be its subset defined as
@@
Then, for a particular sequence $\bm{i}=(i_{1},\ldots,i_{n})$ of the measurement outcomes, a conditional probability of the $n$-th outcome $i_{n}$ conditioned on any sequence $\bm{i}^{\prime}=(i_{1},\ldots,i_{n-1})$ of the other $n-1$ outcomes should be in $\mathcal{Q}$ due to the definition of $\mathcal{Q}$. Let $\bm{l}=(n^{\prime}_{1},\ldots,n^{\prime}_{k})$ with $n^{\prime}_{j}=\chi^{\prime}_{j}(\bm{i}^{\prime})$, where $\bm{\chi}^{\prime}:{\cal X}^{\times(n-1)}\to Z_{n-1}^{k}$ is defined similarly to Eq. (36) for $(n-1)$ tuple, and let $\bm{l}[j]\in Z_{n}^{k}$ be defined as a $k$-tuple that adds one to $j$-th element of $\bm{l}$, i.e., $\bm{l}[j]=(n^{\prime}_{1},\ldots,n^{\prime}_{j-1},n^{\prime}_{j}+1,n^{\prime}_{j+1},\ldots,n^{\prime}_{k})$. By applying the above restriction on the conditional probability to Eq. (127), we have
$\begin{split}\left(\sum_{j=1}^{k}\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}[j])}{|{\cal T}(\bm{l}[j])|}\right)^{-1}\!\!\left(\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}[1])}{|{\cal T}(\bm{l}[1])|},\ldots,\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}[k])}{|{\cal T}(\bm{l}[k])|}\right)\in\mathcal{Q}.\end{split}$ (129)
(Note, if only this condition is satisfied, the following argument holds for any classical probability mass function as well.) This restriction on $\mathrm{P}_{\tilde{\rho}}(\bm{m})$ for each $\bm{m}$ is strict, but is not very convenient for the later analysis. So, we consider a looser version for this that is easier to analyze.
Let ${\cal R}\coloneqq\{(p_{1},\ldots,p_{k})\in{\cal P}:\sum_{j=1}^{k}a_{j}p_{j}\leq b\}$ be a convex subset of ${\cal P}$ with an affine boundary such that ${\cal Q}\subseteq{\cal R}$ (see Fig. 1). We restrict our attention to the case where the complement ${\cal P}\setminus{\cal R}$ of ${\cal R}$ contains only one probability vector that corresponds to having a deterministic outcome, say $\bm{p}_{1}\coloneqq(1,0,\ldots,0)$. This may lead to a looser bound when applied to a specific problem or limit the applicability of the obtained bound, but as can be shown later, it still has a wide range of applications. The coefficients $\{a_{j}\}_{j=1}^{k}$ that defines the set ${\cal R}$ in this case should satisfy $a_{1}>b$, $a_{j}\leq b$ for $j=2,\ldots,k$. By redefining $b_{j}=-(a_{j}-b)/(a_{1}-b)(\geq 0)$ for $j=2,\ldots,k$, we have
${\cal R}$ =$=\left\{(p_{1},\ldots,p_{k})\in{\cal P}:p_{1}-\biggl{(}\sum_{j=2}^{k}b_{j}p_{j}\biggr{)}\leq 0\right\}$ (130)
@@
where $\bm{u}\coloneqq(1,-b_{2},\ldots,-b_{k})$ is as defined in the theorem statement. In this way, we obtain the expression of ${\cal R}$ in Eq. (57). In other words, Eq. (57) is defined to satisfy the condition on the complement ${\cal P}\setminus{\cal R}$ stated above. We also define ${\cal R}^{\prime}$ as in the theorem statement, i.e., ${\cal R}^{\prime}\coloneqq\{\bm{p}\in{\cal P}:\bm{u}\cdot\bm{p}\leq\frac{\sqrt{2}\|\bm{u}-(\bm{v}\cdot\bm{u})\bm{v}\|}{n}\}$, where $\bm{v}\coloneqq(1,\ldots,1)/\sqrt{k}$ is as defined in the theorem statement. Since ${\cal Q}\subseteq{\cal R}$ by assumption, we have from Eq. (129) that
$\begin{split}\left(\sum_{j=1}^{k}\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}[j])}{|{\cal T}(\bm{l}[j])|}\right)^{-1}\!\!\left(\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}[1])}{|{\cal T}(\bm{l}[1])|},\ldots,\frac{\mathrm{P}_{\tilde{\rho}}(\bm{l}[k])}{|{\cal T}(\bm{l}[k])|}\right)\in\mathcal{R}.\end{split}$ (132)
Then, with this loosened constraint, we will find an upper bound on $\mathrm{P}_{\tilde{\rho}}(\bm{m})$ for an arbitrary frequency distribution $\bm{m}=(n_{1},\ldots,n_{k})\in Z_{n}^{k}$ of the outcomes.
As a first step, we will show that given an arbitrary probability vector $\bm{p}\in{\cal R}$, any probability vector $\bm{q}\in{\cal P}$ with $\|\bm{p}-\bm{q}\|\leq\sqrt{2}/n$ is contained in ${\cal R}^{\prime}$. From the Cauchy-Schwarz inequality, we have
$\bm{u}\cdot(\bm{q}-\bm{p})$ $=(\bm{u}-(\bm{v}\cdot\bm{u})\bm{v})\cdot(\bm{q}-\bm{p})$ (133)
$\leq\|\bm{u}-(\bm{v}\cdot\bm{u})\bm{v}\|\|\bm{q}-\bm{p}\|$ (134)
$\leq\frac{\sqrt{2}\|\bm{u}-(\bm{v}\cdot\bm{u})\bm{v}\|}{n},$ (135)
where we used $\bm{v}\cdot\bm{p}=\sqrt{k}=\bm{v}\cdot\bm{q}$ in the first equality. (Note that $\bm{u}-(\bm{v}\cdot\bm{u})\bm{v}$ is a vector projection of $\bm{u}$ onto the affine hull of ${\cal P}$ in $\mathbb{R}^{k}$.) Since $\bm{p}\in{\cal R}$ implies $\bm{u}\cdot\bm{p}\leq 0$, we have from Eq. (135) that
$\bm{u}\cdot\bm{q}\leq\bm{u}\cdot\bm{p}+\frac{\sqrt{2}\|\bm{u}-(\bm{v}\cdot\bm{u})\bm{v}\|}{n}\leq\frac{\sqrt{2}\|\bm{u}-(\bm{v}\cdot\bm{u})\bm{v}\|}{n},$ (136)
which then implies that $\bm{q}\in{\cal R}^{\prime}$ from the definition of ${\cal R}^{\prime}$.
Now, let $\tilde{\cal P}$ be the set of types for $n$ outcomes defined as
$\tilde{\cal P}\coloneqq\{(n_{1},\ldots,n_{k})/n\in{\cal P}:\forall j\in\{1,\ldots,k\},n_{j}\in\mathbb{Z}\},$ (137)
and let $\tilde{\cal R}$ be its subset defined as